This Question: 1 pt 14 of 15 (11 complete) This Quiz: 15 pts possible Assume that when adults with smartphones are randomly selected, 55% use them in meetings or classes. If 14 adult smartphone users are randomly selected, find the probability that fewer than 5 of them use their smartphones in meetings or classes. The probability is (Type an integer or decimal rounded to four decimal places as needed.) Enter your answer in the answer box.

Thank you!!
Answer #1

p = 55% = 0.55, q = 1-p = 0.45

and n14 rl

We need to find P(X < 5)

P(X=x)=n_{C_{x}}(q)^{n-x}(p)^{x}

P(X< 5)=P(X=0)+P(X=1)+P(X=2)+P(X=3)+P(X=4)

  = 14c0(0.15)14-0(0.55) +14cì (0.45) 14-1 (0.55)사 14C2(0.45)14-2 (0.55)2 + 14c, (0.45)14-3(0.55)3 +14c, (0.45)14 (0.55)4 860 591035440430400 + 10.5)

  - 1(0.45)141 +14(0.45)13 (0.55)91(0.45)12(0.55)2 +364(0.45)11 (0.55)3 + 1001 (0.45)10 (0.55)4

0.0426